OpenStudy (anonymous):

Solve -14 = 2/3 t t=

OpenStudy (anonymous):

ok so you take -14 and divide it by 2/3 and you get T.

OpenStudy (anonymous):

so like this ((-14) / 2) / 3 = -2.33333333333

OpenStudy (anonymous):

@iGreen am i correct?

OpenStudy (bookworm14):

\[-14 \div \frac{ 2 }{ 3 } = -21\] t= -21 because -21 * 2/3 - -14 Hope this helps~Bree

OpenStudy (anonymous):

Thanks Bree

OpenStudy (igreen):

Actually, Dividing 2/3 is the same thing as multiplying 3/2. What's \(\dfrac{3}{2} \times -14\)?

OpenStudy (anonymous):

-21

OpenStudy (igreen):

It should be: ((-14) / 2/3) =

OpenStudy (igreen):

Yes, so -21 is your answer.

OpenStudy (igreen):

We never divide fractions, we always flip it and multiply.

OpenStudy (anonymous):

Ok Thanks Green :)

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

^

OpenStudy (bookworm14):

Your very Welcome @BSnowBroGaming , im so happy I could help :D